### v3

# On the multiplicity of weak solutions for a class of coupled quasilinear elliptic systems

Abstract.
We study the existence and regularity of weak solutions to the following quasilinear elliptic system:
$\begin{cases}-\mathrm{div}(A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k})+\dfrac{1}{p_{k}}D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}=g_{k}(x,u)\quad\text{in }\Omega,\\ u_{k}=0\quad\text{on }\partial\Omega,\qquad k=1,\dots,d,\end{cases}$
where $\Omega\subset\mathbb{R}^{N}$ is a bounded domain with $N\geq 2$, $\boldsymbol{p}=(p_{1},\dots,p_{d})$, $p_{k}>1$. Using tools from nonsmooth critical point theory, we prove the existence of infinitely many weak solutions in $W_{0}^{1,\boldsymbol{p}}(\Omega)\cap L^{\infty}(\Omega;\mathbb{R}^{d})$, where $W_{0}^{1,\boldsymbol{p}}(\Omega)=W_{0}^{1,p_{1}}(\Omega)\times\dots\times W_{0}^{1,p_{d}}(\Omega)$.
Keywords: Subcritical nonlinearities, gradient elliptic systems, Dirichlet boundary conditions, quasilinear elliptic equations, nonsmooth critical point theory.
2020 MSC: 35A01, 35A15, 35J05, 35J20, 35J25.
1. Introduction
Let $\Omega\subset\mathbb{R}^{N}$ be a bounded domain with $N\geq 2$, $p_{1},\dots,p_{d}>1$, $\boldsymbol{p}=(p_{1},\dots,p_{d})$, and
$W_{0}^{1,\boldsymbol{p}}(\Omega):=W_{0}^{1,p_{1}}(\Omega)\times\dots\times W_{0}^{1,p_{d}}(\Omega).$
We consider the functional $f:W_{0}^{1,\boldsymbol{p}}(\Omega)\to\mathbb{R}$, defined as follows:
$f(u)=\sum_{k=1}^{d}\frac{1}{p_{k}}\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}-\int_{\Omega}G(x,u).$
For every $k=1,\dots,d$, we assume that
  • $A_{k}:\Omega\times\mathbb{R}\to\mathbb{R}$ is a $C^{1}$-Carathéodory function, i.e., $A_{k}(\cdot,s)$ is measurable for every $s\in\mathbb{R}$, and $A_{k}(x,\cdot)$ is $C^{1}$ for a.e. $x\in\Omega$.
  • There exist constants $C_{0},\nu>0$ such that
    ($a.1$) $|A_{k}(x,s)|,|D_{s}A_{k}(x,s)|\leq C_{0},$
    ($a.2$) $A_{k}(x,s)\geq\nu>0,$
    for a.e. $x\in\Omega$, for every $s\in\mathbb{R}$,
  • $G:\Omega\times\mathbb{R}^{d}\to\mathbb{R}$ is a $C^{1}$-Carathéodory function such that $G_{s_{k}}(x,0):=D_{s_{k}}G(x,0)=0$ and $G(x,0)=0$,
  • there exists $C>0$ such that
    ($g.1$) $\left|G_{s_{k}}(x,s)\right|\leq C\left(1+|s_{k}|^{q_{k}-1}+\sum_{j\neq k}|s_{j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}\right),\quad\text{and $q_{k}\in(p_{k},p_{k}^{*})$},$
    for a.e. $x\in\Omega$ and for every $s\in\mathbb{R}^{d}$, where
    $p_{k}^{*}=\begin{cases}\frac{Np_{k}}{N-p_{k}},&\text{if }N>p_{k},\\ \infty,&\text{otherwise}.\end{cases}$
  • There exist constants $R>0$, $\mu>\overline{p}:=\max\{p_{1},\dots,p_{d}\}$, and $\gamma\in(0,\mu-\overline{p})$ such that
    ($g.2$) $|s|\geq R\implies 0<\mu G(x,s)\leq\sum_{k=1}^{d}s_{k}G_{s_{k}}(x,s),$
    ($a.3$) $|s|\geq R\implies 0\leq sD_{s}A_{k}(x,s)\leq\gamma A_{k}(x,s)$
    for a.e. $x\in\Omega$.
  • We denote $g_{k}(x,s):=G_{s_{k}}(x,s)$.
In these assumptions, the energy functional is continuous in $W_{0}^{1,\boldsymbol{p}}(\Omega)$ (as in Theorem 2.7 below), but it is Gâteaux differentiable only along the directions $v\in W_{0}^{1,\boldsymbol{p}}(\Omega)\cap L^{\infty}(\Omega;\mathbb{R}^{d})$ and
$\langle f^{\prime}(u),v\rangle=\sum_{k=1}^{d}\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla v_{k}$ $+\sum_{k=1}^{d}\frac{1}{p_{k}}\int_{\Omega}D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}v_{k}$
$-\sum_{k=1}^{d}\int_{\Omega}g_{k}(x,u)v_{k}.$
Hence, a function $u\in W_{0}^{1,\boldsymbol{p}}(\Omega)$ such that $\langle f^{\prime}(u),v\rangle=0$ for every $v\in C_{c}^{\infty}(\Omega;\mathbb{R}^{d})$ is a weak solution in $\mathcal{D}^{\prime}(\Omega)$ for
$\begin{cases}-\mathrm{div}(A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k})+\dfrac{1}{p_{k}}D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}=g_{k}(x,u)\quad\text{in }\Omega,\\ u_{k}=0\quad\text{on }\partial\Omega,\qquad k=1,\dots,d.\end{cases}$
Namely:
(1.1) $\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla v_{k}+\frac{1}{p_{k}}\int_{\Omega}D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}v_{k}=\int_{\Omega}g_{k}(x,u)v_{k},$
for every $v\in C_{c}^{\infty}(\Omega;\mathbb{R}^{d})$ and $k=1,\dots,d$.
The main aim of this paper is to study the existence of critical points of $f$ in the framework of nonsmooth critical point theory [12, 16], following the approach developed in [10, 11, 12, 13, 19].
In particular, the existence of infinitely many solutions to the Dirichlet problem
$\operatorname{div}(A(x,u)\nabla u)+\frac{1}{2}\,D_{s}A(x,u)\nabla u\cdot\nabla u=g(x,u)\quad\text{in }\Omega,\qquad u=0\quad\text{on }\partial\Omega,$
where $A(x,s)$ is a matrix with $C^{1}$-Carathéodory coefficients.
A recent extension of this approach to the corresponding Neumann problem
$\operatorname{div}(A(x,u)\nabla u)+\frac{1}{2}\,D_{s}A(x,u)\nabla u\cdot\nabla u=g(x,u)\quad\text{in }\Omega,\qquad A(x,u)\nabla u\cdot\eta=0\quad\text{on }\partial\Omega,$
is presented in [13].
Moreover, a broader class of quasilinear equations of the form
$\operatorname{div}(a(x,u,\nabla u))+b(x,u,\nabla u)=g(x,u)\quad\text{in }\Omega,\qquad u\in W_{0}^{1,p}(\Omega;\mathbb{R}),$
has been studied in [19], where the same nonsmooth variational framework is employed to obtain multiplicity results. We also mention the related works [2, 6, 7], which deal with the same type of quasilinear problems but consider the functional on the subspace $W_{0}^{1,p}(\Omega)\cap L^{\infty}(\Omega)$, where it becomes differentiable. However, in this way the treatment of the Palais-Smale condition becomes more involved.
Several classes of elliptic systems have been investigated in the literature through different analytical approaches. Variational systems—namely those that arise as Euler–Lagrange equations of a suitable functional—can be studied by means of smooth or nonsmooth critical point theory, depending on the regularity of the associated energy functional. Classical examples are given by gradient systems of the form
$\Delta u_{1}=g_{1}(x,u_{1},u_{2}),\qquad-\Delta u_{2}=g_{2}(x,u_{1},u_{2}),$
for which existence and multiplicity of solutions can be obtained through minimization techniques, Mountain Pass arguments, or linking constructions; see, for instance, [15].
The existence of solutions for systems involving the $p$-Laplacian, such as
$\Delta_{p_{1}}u_{1}=g_{1}(x,u),\qquad-\Delta_{p_{2}}u_{2}=g_{2}(x,u),\qquad u=(u_{1},u_{2})\in W_{0}^{1,p_{1}}(\Omega)\times W_{0}^{1,p_{2}}(\Omega),$
was studied in [3] under the growth condition ($g.1$). These operators appear in models of non-Newtonian fluids, nonlinear elasticity, and glaciology, where the diffusion operator is of $p$-Laplacian type, and the coupling reflects the interaction between different physical or mechanical components.
In the nonsmooth setting, results for quasilinear systems have been established in [1, 20], where the authors consider nonlinear systems of the type
$\operatorname{div}(A_{k}(x,u)\nabla u_{k})+\frac{1}{2}\sum_{i=1}^{d}D_{s_{k}}A_{i}(x,u)\,\nabla u_{i}\cdot\nabla u_{i}=g_{k}(x,u),\qquad u=(u_{1},\dots,u_{d})\in W_{0}^{1,2}(\Omega;\mathbb{R}^{d}),$
and prove the existence of multiple bounded solutions within the framework of nonsmooth critical point theory.
More recently, in [8, 9], the authors consider the problem (1) with $d=2$, adopting a different variational framework and assuming a stronger growth condition on the nonlinearities $g_{1}(x,u,v),g_{2}(x,u,v)$.
In the present paper, we assume instead the weaker growth condition ($g.1$), namely the one used in the regular case [3], and we allow $d\geq 2$.
Working directly in the Banach space $W_{0}^{1,p}(\Omega)$, we rely on the nonsmooth critical point theory developed in [12, 16], where the notion of weak slope is introduced for continuous functionals, extending the classical derivative norm $\|f^{\prime}(u)\|$ and providing natural definitions of critical points, Palais–Smale sequences, and the Palais–Smale condition for possibly non-differentiable functionals.
By applying this framework, we prove the existence of infinitely many critical points for $f$ through the nonsmooth version of the Equivariant Mountain Pass Theorem.
Finally, we show that the corresponding weak solutions are bounded, as a consequence of a regularity result based on [22].
Our main result is the following:
Theorem 1.1.
Assume that hypotheses (a.1a.1)-($a.3$), (g.1g.1)-($g.2$) are satisfied and that
$A_{k}(x,-s)=A_{k}(x,s),\qquad G(x,-s)=G(x,s).$
Then, there exists a sequence of weak solutions $\{u_{h}\}\subset W_{0}^{1,\boldsymbol{p}}(\Omega)\cap L^{\infty}(\Omega;\mathbb{R}^{d})$ of (1), such that $f(u_{h})\to+\infty$ as $h\to+\infty$.
2. Preliminaries
2.1. Weak slope
We recall some results about the critical point theory of continuous functionals, developed in [12]. In this setting, we consider $X$ a metric space and $f:X\to\mathbb{R}$ a continuous functional.
Definition 2.1.
Let $X$ be a metric space and let $f:X\to\mathds{R}$ be a continuous function. We consider $\sigma\geq 0$ such that there exist $\delta>0$ and a continuous map $\mathscr{H}:B_{\delta}(u)\times[0,\delta]\to X$ such that
(2.1) $d(\mathscr{H}(v,t),v)\leq t,$
(2.2) $f(\mathscr{H}(v,t))\leq f(v)-\sigma t.$
We define
[…]
as the weak slope of $f$ at $u$.
Theorem 2.2 ([12, Theorem 1.1.2]).
Let $E$ be a normed space and $X\subset E$ an open subset. Let us fix $u\in X$ and $v\in E$ with $\|v\|=1$. For each $w\in X$ we define
$\overline{D}_{+}f(w)[v]:=\limsup_{t\to 0^{+}}\frac{f(w+tv)-f(w)}{t}.$
Then $|df|(u)\geq-\limsup_{w\to u}\overline{D}_{+}f(w)[v]$.
Definition 2.3.
Let $X$ be a metric space and let $f:X\to\mathds{R}$ be continuous. We say that $u\in X$ is a (lower) critical point if $|df|(u)=0$. A (lower) critical point is said at the level $c\in\mathbb{R}$ if it is also true that $f(u)=c$.
Definition 2.4.
Let $X$ be a metric space and let $f:X\to\mathds{R}$ be continuous. A sequence $\{u_{n}\}\subset X$ is a $(PS)_{c}$-sequence if
(2.3) $f(u_{n})\to c,$
(2.4) $|df|(u_{n})\to 0.$
Furthermore, we will say that $f$ satisfies the $(PS)_{c}$-condition if every $(PS)_{c}$-sequence admits a convergent subsequence in $X$. If the $(PS)_{c}$-condition holds for every $c\in\mathbb{R}$ we will simply write $(PS)$-condition.
Theorem 2.5 (Equivariant Mountain Pass, [12, Theorem 1.3.3]).
Let $X$ be a Banach space and let $f:X\to\mathbb{R}$ be a continuous even functional. Suppose that
  • $\exists\ \rho>0,\alpha>f(0)$ and a subspace $X_{2}\subset X$ of finite codimension such that
    $f\geq\alpha\ on\ \partial B_{\rho}\cap X_{2}$,
  • for every subspace $X_{1}^{(k)}$ of dimension $k$, there exists $R=R^{(k)}>0$ such that $f\leq f(0)$ in $B_{R}^{c}\cap X_{1}^{(k)}$.
If $f$ satisfies the $(PS)_{c}$-condition for every $c\geq\alpha$, then there exists a divergent sequence of critical values, namely there exists a sequence $\{u_{n}\}\subset X$ such that $c_{n}:=f(u_{n})\to+\infty$.
2.2. Properties of energy functional
In this subsection, and in the rest of the paper, $f$ is the energy functional
$f(u)=\sum_{k=1}^{d}\frac{1}{p_{k}}\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}-\int_{\Omega}G(x,u).$
We denote the dual of the space $W_{0}^{1,\boldsymbol{p}}$ by $W^{-1,\boldsymbol{p}^{\prime}}$, namely:
$W^{-1,\boldsymbol{p}^{\prime}}:=W^{-1,p_{1}^{\prime}}\times\dots\times W^{-1,p_{d}^{\prime}},$
where $p_{k}^{\prime}$ is the conjugate exponent of $p_{k}$.
Definition 2.6.
A sequence $\{u_{n}\}\subset W_{0}^{1,\boldsymbol{p}}(\Omega)$ is said a Concrete Palais-Smale at level $c$ for $f$, $(CPS)_{c}$-sequence, if
  • $f(u_{n})\to c$ in $\mathbb{R}$.
  • $\text{div}(A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}-2}\nabla u_{n,k})+\frac{1}{p_{k}}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}-g_{k}(x,u_{n})\in W^{-1,p_{k}^{\prime}}(\Omega)$, eventually as $n\to+\infty$, for every $k=1,\dots,d$.
  • $\text{div}(A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}-2}\nabla u_{n,k})+\frac{1}{p_{k}}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}-g_{k}(x,u_{n})\to 0$ strongly in $W^{-1,p_{k}^{\prime}}(\Omega)$ for every $k=1,\dots,d.$
We will say that $f$ satisfies the $(CPS)_{c}$-condition if every $(CPS)_{c}$-sequence admits a convergent subsequence in $W_{0}^{1,\boldsymbol{p}}(\Omega)$.
Theorem 2.7.
The functional $f:W_{0}^{1,\boldsymbol{p}}(\Omega)\to\mathbb{R}$ is continuous and for every $u\in W_{0}^{1,\boldsymbol{p}}(\Omega)$ we have that
$|df|(u)\geq\sup_{\begin{subarray}{c}\varphi\in C_{c}^{\infty}(\Omega;\mathbb{R}^{d})\\ \|\varphi\|_{W_{0}^{1,\boldsymbol{p}}(\Omega)}\leq 1\end{subarray}}\sum_{k=1}^{d}\biggl[\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla\varphi_{k}$ $+\frac{1}{p_{k}}\int_{\Omega}D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\varphi_{k}$
$-\int_{\Omega}g_{k}(x,u)\varphi_{k}\biggr].$
Proof.
The continuity of the functional follows from the fact that $A_{k}(x,u_{k})$ are continuous in $u_{k}$ and by the Lebesgue’s Theorem. Let, now, $\varphi\in C_{c}^{\infty}(\Omega;\mathbb{R}^{d})$ . Then by Lebesgue’s Theorem, the quantity
$\langle f^{\prime}(u),\varphi\rangle:$ $=\lim_{t\to 0}\frac{f(u+t\varphi)-f(u)}{t}$
$=\sum_{k=1}^{d}\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla\varphi_{k}+\sum_{k=1}^{d}\frac{1}{p_{k}}\int_{\Omega}D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\varphi_{k}$
$\quad-\sum_{k=1}^{d}\int_{\Omega}g_{k}(x,u)\varphi_{k}$
is well defined. Furthermore, from Theorem 2.2, we have that
$\langle f^{\prime}(u),\varphi\rangle=\limsup_{\begin{subarray}{c}w\to u\\ \ t\to 0^{+}\end{subarray}}\frac{f(w+t\varphi)-f(w)}{t}\geq-|df|(u),\qquad\text{if $\|\varphi\|=1$.}$
In particular, $|df|(u)\geq\langle f^{\prime}(u),-\varphi\rangle$ and, by arbitrariness of $\varphi$, the thesis follows. ∎
Corollary 2.8.
The following facts hold:
  1. $(i)$
    if $u$ is a lower critical point for $f$, then $u$ is a weak solution of the problem (1),
  2. $(ii)$
    every $(PS)_{c}$-sequence is also a $(CPS)_{c}$-sequence,
  3. $(iii)$
[…]
We conclude this section with a useful Brezis-Browder type result. This result follows as a consequence of [4]. However, for completeness, we provide an explicit proof in our setting.
Theorem 2.9.
Let $\omega\in W^{-1,\boldsymbol{p}^{\prime}}(\Omega)$ and $u\in W_{0}^{1,\boldsymbol{p}}(\Omega)$ such that
[…]
Let $v\in W_{0}^{1,\boldsymbol{p}}(\Omega)$ such that $\left(D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}v_{k}\right)^{-}\in L^{1}(\Omega)$ for every $k=1,\dots,d$. Then, one has that
$D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}v_{k}\in L^{1}(\Omega)\ \ \text{and}$
$\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla v_{k}+\frac{1}{p_{k}}\int_{\Omega}D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}v_{k}=\langle\omega_{k},v_{k}\rangle,$
for every $k=1,\dots,d$.
Proof.
The case $v\in W_{0}^{1,\boldsymbol{p}}(\Omega)\cap L^{\infty}(\Omega;\mathbb{R}^{d})$ follows by regularizing $v$ via convolution (or simply using Riesz’s Theorem). Let $\{\tilde{v}_{n}\}\subset C_{c}^{\infty}(\Omega;\mathbb{R}^{d})$ such that $\tilde{v}_{n}\to v$ in $W_{0}^{1,\boldsymbol{p}}(\Omega)$ and we consider, as in [5], $v_{n,k}(x)=v_{k}(x)\cdot\lambda_{n,k}(x)$, with $\lambda_{n,k}(x)\in[0,1]^{d}$ and
$v_{n,k}=v_{k}\bigg(|v_{k}|^{p_{k}}+\frac{1}{n^{p_{k}}}\bigg)^{-\frac{1}{p_{k}}}\min\bigg\{\bigg(|v_{k}|^{p_{k}}+\frac{1}{n^{p_{k}}}\bigg)^{\frac{1}{p_{k}}}-\frac{1}{n},\bigg(|\tilde{v}_{n,k}|^{p_{k}}+\frac{1}{n^{p_{k}}}\bigg)^{\frac{1}{p_{k}}}-\frac{1}{n}\bigg\}\in W_{0}^{1,p_{k}}\cap L^{\infty}.$
We point out that $v_{n,k}\to v_{k}$ in $W_{0}^{1,p_{k}}(\Omega)$, $|v_{n,k}|\leq|v_{k}|$ a.e. in $\Omega$ with $k=1,\dots,d$. Let us define
$T_{k}(x):=D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}},\qquad\text{for a.e. $x\in\Omega$.}$
Hence,
$T_{k}(x)v_{n,k}(x)=\lambda_{n,k}(x)D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}v_{k}=\lambda_{n,k}(x)T_{k}(x)v_{k}(x),$
and $0\leq\lambda_{n,k}\leq 1.$ Then
$T_{k}v_{n,k}=\lambda_{n,k}T_{k}v_{k}\geq-\lambda_{n,k}\left[D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\right]^{-}\geq-\left[D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}v_{k}\right]^{-}.$
According to Fatou’s Lemma:
$\int_{\Omega}T_{k}v_{k}+\int_{\Omega}\left[D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}v_{k}\right]^{-}\leq\liminf_{n\to+\infty}\left\{\int_{\Omega}T_{k}v_{n,k}+\int_{\Omega}\left[D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}v_{k}\right]^{-}\right\}.$
By (2.5) we obtain
$\int_{\Omega}D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}v_{k}$ $\leq\liminf_{n\to+\infty}\int_{\Omega}D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}v_{n,k}$
$=p_{k}\liminf_{n\to+\infty}\bigg\{\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla v_{n,k}-\langle\omega_{k},v_{n,k}\rangle\bigg\}$
$\leq C_{k},$
for some positive constant $C_{k}>0$. Thus, $D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}v_{k}\in L^{1}(\Omega)$. Testing (2.5) with $v_{n,k}$, we can pass to the limit according to Lebesgue’s Theorem and obtain:
$\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla v_{k}+\frac{1}{p_{k}}\int_{\Omega}D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}v_{k}=\langle\omega_{k},v_{k}\rangle.$
Remark 2.10.
Notice that
$\left(D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\right)v_{k}\in L^{1}(\Omega),\quad k=1,\dots,d,$
for every $v\in L^{\infty}(\Omega;\mathbb{R}^{d})$ and $u\in W_{0}^{1,\boldsymbol{p}}(\Omega)$. Hence, Theorem 2.9 implies that $v\in W_{0}^{1,\boldsymbol{p}}(\Omega)\cap L^{\infty}(\Omega;\mathbb{R}^{d})$ is an admissible test function for (1).
3. Regularity result
Now, we investigate the $L^{\infty}$-regularity of the weak solutions.
Theorem 3.1.
Let $u=(u_{1},\dots,u_{d})\in W_{0}^{1,\boldsymbol{p}}(\Omega)$ be a weak solution of (1). Then $u_{1},\dots,u_{d}\in L^{\infty}(\Omega)$.
Proof.
We adapt the procedure in [22], see also [14] and [17]. For every $\gamma,t,L$ such that $L>R+1$, we define
$\tilde{h}_{L,\gamma}(s)=\begin{cases}|s|^{\gamma-1}s,&|s|\leq L,\\ \gamma L^{\gamma-1}s-\operatorname{sign}(s)(1-\gamma)L^{\gamma},&|s|>L,\end{cases}$
and $h_{L,\gamma}(s)$ the function such that
  • $h_{L,\gamma}\equiv 0$ for $|s|\leq R$;
  • $h_{L,\gamma}$ is the line which joins $(R,0)$ with the point $(R+1,(R+1)^{\gamma})$ for $s\in(R,R+1]$ and it is the line which joins (R,0)(-R,0) with $(-R-1,-(R+1)^{\gamma})$ for $s\in[-R-1,-R)$;
  • $h_{L,\gamma}\equiv\tilde{h}_{L,\gamma}$ for $|s|>R+1$.
Since $h_{L,\gamma}$ is differentiable a.e. in $\mathbb{R}$ and its derivative is bounded, we can also define:
$\Phi_{L,t,\gamma}(s)=\int_{0}^{s}|h_{L,\gamma}^{\prime}(r)|^{\frac{t}{\gamma}}\,dr.$
Notice that for every $|s|>L$ we have $h_{L,\gamma}^{\prime}\equiv\gamma L^{\gamma-1}$ and this implies that $\Phi_{L,t,\gamma}(s)=c_{1}(L,\gamma,t)s$. Thus, if $|u_{k}|>L$ we have
$D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\Phi_{L,t,\gamma}(u_{k})\geq 0$
by ($a.3$) and, in particular, $D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\Phi_{L,t,\gamma}(u_{k})]^{-}\in L^{1}(\Omega)$. According to Theorem 2.9, $\Phi_{L,t,\gamma}$ is an admissible test function. Now, we start testing (1.1) with $\Phi_{L,\gamma p_{k},\gamma}(u_{k})e_{k}$ where $\{e_{1},\dots,e_{d}\}$ is the canonical basis of $\mathbb{R}^{d}$.
From ($a.3$), we obtain that
$\frac{1}{p_{k}}\int_{\Omega}D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\Phi_{L,\gamma p_{k},\gamma}(u_{k})\geq 0.$
Indeed, $h_{L,\gamma}^{\prime}(s)$ is a constant for $|s|\leq R+1$ and for $|s|>L$, which implies that $\Phi_{L,\gamma p_{k},\gamma}=\theta s$ for some $\theta\geq 0$ and for every $s\in\mathbb{R}$ such that either $|s|\leq R+1$ or $|s|>L$. Otherwise, $h_{L,\gamma}^{\prime}(s)=\gamma|s|^{\gamma-1}$ and $\Phi_{L,\gamma p_{k},\gamma}(s)=s|s|^{\gamma-1}$. Since
$|s|\geq R\implies sD_{s}A_{k}(x,s)\geq 0,$
we obtain that $D_{s}A_{k}(x,s)\Phi_{L,\gamma p_{k},\gamma}(s)\geq 0$ and the claim follows. Furthermore:
$\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla\Phi_{L,\gamma p_{k},\gamma}(u_{k})=\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}|h_{L,\gamma}^{\prime}(u_{k})|^{p_{k}}.$
Since $h_{L,\gamma}(u_{k})\in W_{0}^{1,p_{k}}(\Omega)$, the Sobolev embedding and (1.1) imply that
$\left(\int_{\Omega}|h_{L,\gamma}(u_{k})|^{q_{k}}\right)^{\frac{p_{k}}{q_{k}}}$ $\leq c\int_{\Omega}|\nabla h_{L,\gamma}(u_{k})|^{p_{k}}=\int_{\Omega}|\nabla u_{k}|^{p_{k}}\cdot|h_{L,\gamma}^{\prime}(u_{k})|^{p_{k}}$
$\leq\frac{1}{\nu}\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}|h_{L,\gamma}^{\prime}(u_{k})|^{p_{k}}$
$\leq C_{1}\int_{\Omega}\left(1+|u_{k}|^{q_{k}-1}+\sum_{j\neq k}|u_{j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}\right)|\Phi_{L,\gamma p_{k},\gamma}(u_{k})|.$
Now, we can follow equations (2.7)-(2.10) in [22].
Notice that the estimates [22, (2.3) and (2.4)] hold for $|s|\geq R+1$.
Let $\Omega_{\sigma,w}=\{x\in\Omega\ :\ |w(x)|>\sigma\}$, for some $\sigma>R+1$. The estimate [22, eq (2.7)] becomes:
$\int_{\{|u_{k}|\geq R+1\}}(|u_{k}|^{q_{k}-1}+1)|\Phi_{L,\gamma p_{k},\gamma}(u_{k})|\leq$ $C_{2}\sigma^{q_{k}-1}\left(\int_{\Omega}|h_{L,\gamma}(u_{k})|^{q_{k}}\right)^{\frac{p_{k}}{q_{k}}\frac{\gamma p_{k}+1-p_{k}}{\gamma p_{k}}}$
$+C_{2}\|u_{k}\|_{L^{q_{k}}(\Omega_{\sigma,u_{k}})}^{q_{k}-p_{k}}\left(\int_{\Omega}|h_{L,\gamma}(u_{k})|^{q_{k}}\right)^{\frac{p_{k}}{q_{k}}}.$
We consider the coupling term:
$\int_{\Omega}|u_{j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}|\Phi_{L,\gamma p_{k},\gamma}(u_{k})|$ $\leq\sigma^{q_{j}\frac{q_{k}-1}{q_{k}}}\int_{\Omega}|\Phi_{L,\gamma p_{k},\gamma}(u_{k})|+\int_{\Omega_{\sigma,u_{j}}}|u_{j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}|\Phi_{L,\gamma p_{k},\gamma}(u_{k})|$
$\leq c_{1}(\sigma,R)+\sigma^{q_{j}\frac{q_{k}-1}{q_{k}}}\int_{\{|u_{k}|\geq R+1\}}\Phi_{L,\gamma p_{k},\gamma}(u_{k})|$
$+\int_{\Omega_{\sigma,u_{j}}}|u_{j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}|\Phi_{L,\gamma p_{k},\gamma}(u_{k})|=:c_{1}(\sigma,R)+I_{j,k}.$
The integral $I_{j,k}$ can be treated as in [22, pages 5-6]:
$\sum_{j\neq k}I_{j,k}\leq$ $\sum_{j\neq k}\Bigg\{C_{3}\sigma^{q_{j}\frac{q_{k}-1}{q_{k}}}\left(\int_{\Omega}|h_{L,\gamma}(u_{k})|^{q_{k}}\right)^{\frac{p_{k}}{q_{k}}\frac{\gamma p_{k}+1-p_{k}}{\gamma p_{k}}}$
$+C_{3}\|u_{j}\|_{L^{q_{j}}(\Omega_{\sigma,u_{j}})}^{\frac{q_{j}}{q_{k}}(q_{k}-p_{k})}\left[\left(\int_{\Omega}|h_{L,\gamma}(u_{k})|^{q_{k}}\right)^{\frac{p_{k}}{q_{k}}}+\left(\int_{\Omega}\left|h_{L_{1},\gamma}(u_{j})\right|^{q_{j}}\right)^{\frac{p_{k}}{q_{k}}}\right]\Bigg\},$
where $L_{1}:=L^{\frac{q_{k}}{q_{j}}}$. Thus, [22, eq (2.9)] takes the form:
$\int_{\Omega}|h_{L,\gamma}(u_{k})|^{q_{k}}$ $\leq C_{4}+C_{4}\left(\sigma^{q_{k}-1}+\sum_{j\neq k}\sigma^{\frac{q_{j}}{q_{k}}(q_{k}-1)}\right)^{\frac{\gamma q_{k}}{p_{k}-1}}+C_{4}\sum_{j\neq k}\|u_{j}\|_{L^{q_{j}}(\Omega_{\sigma,u_{j}})}^{\frac{q_{j}}{p_{k}}(q_{k}-p_{k})}\int_{\Omega}|h_{L_{1},\gamma}(u_{j})|^{q_{j}}.$
According to [22, Lemma 2.1], we can choose $\sigma\geq\sigma_{1}$ and $C_{5}>0$ independent on $L$ such that:
$\int_{\Omega}|h_{L,\gamma}(u_{j})|^{q_{j}}\leq C_{5}.$
Passing to the limit as $L\to+\infty$ and since $\gamma>1$ is an arbitrary number, we have that $u=(u_{1},\dots,u_{d})\in L^{t}(\Omega;\mathbb{R}^{d})$ for every $t\in(1,\infty)$. In particular, there exists $m>\max_{1\leq k\leq d}\frac{N}{p_{k}}$ such that $g_{k}(x,u)\in L^{m}(\Omega;\mathbb{R})$ for every $k=1,\dots,d$. We define
$\xi(s)=\begin{cases}s+R&\text{if $s<-R$}\\ 0&\text{if $|s|\leq R$}\\ s-R&\text{if $s>R$}.\end{cases}$
Thus,
$\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla\xi(u_{k})+\frac{1}{p_{k}}\int_{\Omega}D_{s}A(x,u_{k})|\nabla u_{k}|^{p_{k}}\xi(u_{k})=\int_{\Omega}g_{k}(x,u)\xi(u_{k}).$
Moreover
$\int_{\Omega}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla\xi(u_{k})=\int_{\{|u_{k}|\geq R\}}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}-2}\nabla u_{k}\cdot\nabla u_{k},$
$\int_{\Omega}D_{s}A_{k}(x,u_{k})|\nabla u_{k}|^{p_{k}}\xi(u_{k})\geq 0.$
Therefore,
$\nu\left(\int_{\Omega}|\xi(u_{k})|^{p_{k}^{*}}\right)^{\frac{p_{k}}{p_{k}^{*}}}$ $\leq\nu C\int_{\{|u_{k}|>R\}}|\nabla u_{k}|^{p_{k}}\leq\nu C\int_{\Omega}|\nabla\xi(u_{k})|^{p_{k}}\leq C\int_{\Omega}g_{k}(x,u)\xi(u_{k})$
$\leq C\left(\int_{\{|u_{k}|>R\}}|g_{k}(x,u)|^{(p_{k}^{*})^{\prime}}\right)^{\frac{1}{(p_{k}^{*})^{\prime}}}\cdot\left(\int_{\Omega}|\xi(u_{k})|^{p_{k}^{*}}\right)^{\frac{1}{p_{k}^{*}}},$
where $(p_{k}^{*})^{\prime}$ is such that
$\frac{1}{p_{k}^{*}}+\frac{1}{(p_{k}^{*})^{\prime}}=1.$
Hence,
$\nu\left(\int_{\Omega}|\xi(u_{k})|^{p_{k}^{*}}\right)^{\frac{p_{k}-1}{p_{k}^{*}}}$ $\leq C\left(\int_{\{|u|>R\}}|g_{k}(x,u)|^{(p_{k}^{*})^{\prime}}\right)^{\frac{1}{(p_{k}^{*})^{\prime}}}$
$\leq C\left(\int_{\{|u|>R\}}|g_{k}(x,u)|^{m}\right)^{\frac{1}{m}}|\{|u_{k}|>R\}|^{\frac{1}{(p_{k}^{*})^{\prime}}-\frac{1}{m}}.$
On the other hand, for every $M>R$ we have that
$\left(\int_{\Omega}|\xi(u_{k})|^{p_{k}^{*}}\right)^{\frac{p_{k}-1}{p_{k}^{*}}}\geq\left(\int_{\{|u_{k}|>M\}}|\xi(u_{k})|^{p_{k}^{*}}\right)^{\frac{p_{k}-1}{p_{k}^{*}}}\geq(M-R)^{p_{k}-1}|\{|u_{k}|>M\}|^{\frac{p_{k}-1}{p_{k}^{*}}}.$
Then,
$|\{|u_{k}|>M\}|\leq\frac{C_{\nu}}{(M-R)^{p_{k}^{*}}}|\{|u_{k}|>R\}|^{\gamma},\quad\gamma=\frac{p_{k}^{*}}{p_{k}-1}\left(\frac{1}{(p_{k}^{*})^{\prime}}-\frac{1}{m}\right)>1.$
According to [21, Lemma 4.1], there exists $K>0$ such that $|\{|u_{k}|>M\}|=0$ for every $M\geq K$ and this implies that $u_{k}\in L^{\infty}(\Omega)$ for every $k=1,\dots,d$.
As a consequence, Theorem 3.1 and Remark 2.10 imply that any weak solution of (1) is an admissible test function.
4. The Concrete Palais-Smale condition
We recall an important compactness result, which follows by [20, Theorem 3.10].
Theorem 4.1.
Let $\{u_{n}\}\subset W_{0}^{1,\boldsymbol{p}}(\Omega)$ be a bounded sequence with $u_{n}=(u_{n,1},\dots,u_{n,d})$ and let $\omega\in W^{-1,\boldsymbol{p}^{\prime}}(\Omega)$ such that
$\int_{\Omega}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}-2}\nabla u_{n,k}\cdot\nabla v_{k}+\frac{1}{p_{k}}\int_{\Omega}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}v_{k}=\langle\omega_{k},v_{k}\rangle,\quad k=1,\dots,d$
for every $v\in C_{c}^{\infty}(\Omega;\mathbb{R}^{d})$. Then $\{u_{n}\}$ admits a strongly convergent subsequence in $W_{0}^{1,\boldsymbol{p}}(\Omega)$.
As consequence, we can state the following:
Theorem 4.2.
For every $c\in\mathbb{R}$ the following facts are equivalent:
  • $(a)$
    $f$ satisfies the $(CPS)_{c}$-condition;
  • $(b)$
    every $(CPS)_{c}$-sequence for $f$ is bounded in $W_{0}^{1,\boldsymbol{p}}(\Omega)$.
Proof.
The implication $(a)\implies(b)$ is a consequence of the strong convergence. Vice versa, let $\{u_{n}\}\subset W_{0}^{1,\boldsymbol{p}}(\Omega)$ be $(CPS)_{c}$-sequence, with $u_{n}=(u_{n,1},\dots,u_{n,d})$, and assume that $u_{n}\to u=(u_{1},\dots,u_{d})$ in $W_{0}^{1,\boldsymbol{p}}(\Omega)$ up to a subsequence. Thus, up to a further subsequence, $u_{n,k}\to u_{k}$ in $L^{t_{k}}(\Omega)$ with $t_{k}\in(p_{k},p_{k}^{*})$ for every $k=1,\dots,d$ and
$\int_{\Omega}g_{k}(x,u_{n})v\to\int_{\Omega}g_{k}(x,u)v,\ \ \text{for every $v_{k}\in W_{0}^{1,p_{k}}(\Omega)$ and $k=1,\dots,d$.}$
According to Theorem 4.1 with $\omega_{n}=(g_{1}(x,u_{n}),\dots,g_{d}(x,u_{n}))$, $\{u_{n}\}$ admits a strongly convergent subsequence. ∎
Proposition 4.3.
Let $c\in\mathbb{R}$ and let $\{u_{n}\}$ be a $(CPS)_{c}$-sequence for the functional $f$ with $u_{n}=(u_{n,1},\dots,u_{n,d})$. Then for every $\rho>0$ and $\varepsilon>0$, there exists $K(\rho,\varepsilon)>0$ such that for every $n\in\mathbb{N}$ and for every $k=1,\dots,d$
$\int_{\Omega_{1}^{\rho}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\leq\varepsilon\int_{\Omega_{2}^{\rho}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K(\rho,\varepsilon),$
$\Omega_{1}^{\rho}=\{x\in\Omega\ :\ |u_{n,k}|\leq\rho\},$ $\Omega_{2}^{\rho}=\{x\in\Omega\ :\ |u_{n,k}|>\rho\}.$
Proof.
We are following [11, Theorem 4.4]. Let $k=1,\dots,d$ and
$\omega_{n,k}:=-\text{div}(A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}-2}\nabla u_{n,k})+\frac{1}{p_{k}}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}-g_{k}(x,u_{n}).$
We have that $\omega_{n,k}\to 0$ in $W^{-1,\boldsymbol{p}^{\prime}}(\Omega)$ as $n\to+\infty$ and for every $k=1,\dots,d$. We use an iterative procedure.
For each $\sigma>0$ and $s=(s_{1},\dots,s_{d})$, we define
$\theta_{1,k}(s)=\begin{cases}s_{k}&|s_{k}|<\sigma\\ -s_{k}+2\sigma&\sigma\leq s_{k}\leq 2\sigma\\ -s_{k}-2\sigma&-2\sigma\leq s_{k}\leq-\sigma\\ 0&|s_{k}|\geq 2\sigma.\end{cases}$
We also set $\theta_{1}=(\theta_{1,1},\dots,\theta_{1,d}):=(\theta_{1,1}(u_{n}),\dots,\theta_{1,d}(u_{n}))$. Then, we have
$\int_{\Omega}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}-2}\nabla u_{n,k}\cdot\nabla\theta_{1,k}+\frac{1}{p_{k}}\int_{\Omega}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\theta_{1,k}$
$\quad=\int_{\Omega}g_{k}(x,u_{n})\theta_{1,k}+\langle\omega_{n,k},\theta_{1,k}\rangle\leq\int_{\Omega}g_{k}(x,u_{n})\theta_{1,k}+\|\omega_{n,k}\|_{W^{-1,p_{k}^{\prime}}}\cdot\|\theta_{1,k}\|_{W_{0}^{1,p_{k}}}.$
According to ($g.1$), we obtain
$\int_{\Omega}g_{k}(x,u_{n})\theta_{1,k}$ $\leq C\left(\int_{\Omega}\theta_{1,k}+\int_{\Omega}|u_{n,k}|^{q_{k}-1}\theta_{1,k}+\sum_{j\neq k}\int_{\Omega}|u_{n,j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}\theta_{1,k}\right).$
Since $|\theta_{1,k}|\leq\sigma$ and $\{u_{n,j}\}$ is uniformly bounded in $L^{q_{j}\frac{q_{k}-1}{q_{k}}}(\Omega)$, we get
$\int_{\Omega}g_{k}(x,u_{n})\theta_{1,k}\leq\tilde{C},\qquad\tilde{C}=\tilde{C}(\sigma).$
From weighted Young’s inequality:
(4.1) $\|\omega_{n,k}\|_{W^{1,p_{k}^{\prime}}}\cdot\|\theta_{1,k}\|_{W_{0}^{1,p_{k}}}\leq\frac{1}{\beta^{p_{k}^{\prime}}p_{k}^{\prime}}\|\omega_{n,k}\|_{W^{-1,p_{k}^{\prime}}}^{p_{k}^{\prime}}+\frac{\beta^{p_{k}}}{p_{k}}\|\theta_{1,k}\|_{W_{0}^{1,p_{k}}}^{p_{k}},$
for every $\beta>0$. On the other hand, from ($a.2$):
$\|\theta_{1,k}\|_{W_{0}^{1,p_{k}}}^{p_{k}}$ $=\int_{\Omega}|\nabla\theta_{1,k}|^{p_{k}}\leq\frac{1}{\nu}\int_{\Omega}A_{k}(x,u_{n,k})|\nabla\theta_{1,k}|^{p_{k}}$
$=\frac{1}{\nu}\int_{\{|u_{n,k}|\leq\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+\frac{1}{\nu}\int_{\{\sigma\leq|u_{n,k}|\leq 2\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}.$
Thus,
$\int_{\{|u_{n,k}|\leq\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}-\int_{\{\sigma\leq|u_{n,k}|<2\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$
$\quad+\frac{1}{p_{k}}\int_{\{|u_{n,k}|\leq\sigma\}}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\theta_{1,k}+\frac{1}{p_{k}}\int_{\{\sigma\leq|u_{n,k}|\leq 2\sigma\}}D_{s}A(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\theta_{1,k}$
$=\int_{\Omega}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}-2}\nabla u_{n,k}\cdot\nabla\theta_{1,k}+\frac{1}{p_{k}}\int_{\Omega}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\theta_{1,k}$
$=\int_{\Omega}g_{k}(x,u_{n})\theta_{1,k}+\langle\omega_{n,k},\theta_{1,k}\rangle\leq\int_{\Omega}g_{k}(x,u_{n})\theta_{1,k}+\|\omega_{n,k}\|_{W^{-1,p_{k}^{\prime}}}\cdot\|\theta_{1,k}\|_{W_{0}^{1,p_{k}}}$
$\leq\tilde{C}(\sigma)+\frac{1}{\beta^{p_{k}^{\prime}}p_{k}^{\prime}}\|\omega_{n,k}\|_{W^{-1,p_{k}^{\prime}}}^{p_{k}^{\prime}}+\frac{\beta^{p_{k}}}{p_{k}}\|\theta_{1,k}\|_{W_{0}^{1,p_{k}}}^{p_{k}}$
$\leq\tilde{C}(\sigma)+\frac{1}{\beta^{p_{k}^{\prime}}p_{k}^{\prime}}\|\omega_{n,k}\|_{W^{-1,p_{k}^{\prime}}}^{p_{k}^{\prime}}+\frac{\beta^{p_{k}}}{\nu p_{k}}\int_{\{|u_{n,k}|\leq\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$
$\quad+\frac{\beta^{p_{k}}}{\nu p_{k}}\int_{\{\sigma\leq|u_{n,k}|\leq 2\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}.$
Now, we define
$M:=\operatorname*{ess\,sup}_{\begin{subarray}{c}x\in\Omega\\ s\in\mathbb{R}\end{subarray}}\frac{|D_{s}A_{k}(x,s)|}{p_{k}A_{k}(x,s)}\leq\frac{C_{0}}{\nu p_{k}},$
where the bound is obtained by ($a.1$) and ($a.2$). Hence,
(4.2) $\frac{1}{p_{k}}|D_{s}A_{k}(x,s)|\cdot|\xi|^{p_{k}}\leq MA_{k}(x,s)|\xi|^{p_{k}},$
for a.e. $x\in\Omega$, $s\in\mathbb{R}$, $\xi\in\mathbb{R}^{N}$ and
$\frac{1}{p_{k}}\int_{\{|u_{n,k}|\leq\sigma\}}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\theta_{1,k}$ $+\frac{1}{p_{k}}\int_{\{\sigma\leq|u_{n,k}|\leq 2\sigma\}}D_{s}A(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\theta_{1,k}$
$\leq M\sigma\int_{\{|u_{n,k}|\leq\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$ $+M\sigma\int_{\{\sigma\leq|u_{n,k}|\leq 2\sigma\}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}.$
Therefore, setting
$I_{r}=\{x\in\Omega:\ |u_{n,k}|\leq r\},\qquad A_{r}=\{x\in\Omega:\ r\leq|u_{n,k}|\leq r+\sigma\},$
we obtain
$\left(1-\sigma M-\frac{\beta^{p_{k}}}{\nu p_{k}}\right)\int_{I_{\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$ $\leq\left(1+\sigma M+\frac{\beta^{p_{k}}}{\nu p_{k}}\right)\int_{A_{\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$
$\quad+K_{2}(\sigma,\beta).$
Choosing $\sigma=\frac{1}{2M}$ and $\beta>0$ such that $\frac{\beta^{p_{k}}}{\nu p_{k}}=\frac{1}{4}$, we obtain
(4.3) $\int_{I_{\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\leq K_{1}\int_{A_{\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K_{2}.$
Similarly, if we consider
$\theta_{2,k}(s)=\begin{cases}0&|s_{k}|\leq\sigma\\ s_{k}-\sigma&\sigma<s_{k}<2\sigma\\ s_{k}+\sigma&-2\sigma<s_{k}<-\sigma\\ -s_{k}+3\sigma&2\sigma\leq s_{k}<3\sigma\\ -s_{k}-3\sigma&-3\sigma\leq s_{k}<-2\sigma\\ 0&|s_{k}|\geq 3\sigma,\end{cases}$
we get
(4.4) $\int_{A_{\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\leq K_{1}^{\prime}\int_{A_{2\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K_{2}^{\prime}.$
Notice that the left hand side is equal to
$\int_{I_{2\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}-\int_{I_{\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}.$
Moreover, by (4.3)-(4.4):
$-\int_{I_{\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$ $\geq-K_{1}\int_{A_{\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}-K_{2}$
$\geq-K_{1}K_{1}^{\prime}\int_{A_{2\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}-K_{1}K_{2}^{\prime}-K_{2}.$
Thus,
$\int_{I_{2\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\leq K_{1}^{\prime\prime}\int_{A_{2\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K_{2}^{\prime\prime}.$
Iterating the procedure, for each $l\geq 1$ we get
(4.5) $\int_{I_{l\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\leq K_{1}^{(l)}\int_{A_{l\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K_{2}^{(l)}.$
Let $l\geq 1,\sigma>0$ such that $l\sigma\geq\max\{\rho,R\}$. Take $\delta\in(0,1)$ and
$\theta_{k}(s)=\begin{cases}0&\text{$|s_{k}|\leq l\sigma$}\\ s_{k}-l\sigma&\text{$l\sigma<s_{k}<(l+1)\sigma$}\\ s_{k}+l\sigma&\text{$-(l+1)\sigma<s_{k}<-l\sigma$}\\ -\delta\,s_{k}+\bigl[1+\delta(l+1)\bigr]\sigma,&(l+1)\sigma\leq s_{k}<(l+1)\sigma+\dfrac{\sigma}{\delta},\\ -\delta\,s_{k}-\bigl[1+\delta(l+1)\bigr]\sigma,&-(l+1)\sigma-\dfrac{\sigma}{\delta}<s_{k}\leq-(l+1)\sigma,\\ 0,&|s_{k}|\geq(l+1)\sigma+\dfrac{\sigma}{\delta}.\end{cases}$
We call $\theta(s)=(\theta_{1}(s),\dots,\theta_{d}(s))$. Hence,
(4.6) $\int_{\Omega}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}-2}\nabla u_{n,k}\cdot\nabla\theta_{k}+\frac{1}{p_{k}}\int_{\Omega}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\theta_{k}$
$\quad=\int_{\Omega}g_{k}(x,u_{n})\theta_{k}+\langle\omega_{n,k},\theta_{k}\rangle\leq\int_{\Omega}g_{k}(x,u_{n})\theta_{k}+\|\omega_{n,k}\|_{W^{-1,p_{k}^{\prime}}}\cdot\|\theta_{k}\|_{W_{0}^{1,p_{k}}}.$
According to ($g.1$)
$\int_{\Omega}g_{k}(x,u_{n})\theta_{k}+\langle\omega_{n,k},\theta_{k}\rangle$ $\leq C\left(\int_{\Omega}\theta_{k}+\int_{\Omega}|u_{n,k}|^{q_{k}-1}\theta_{k}+\sum_{j\neq k}\int_{\Omega}|u_{n,j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}\theta_{k}\right)$
$\quad+\|\omega_{n,k}\|_{W^{-1,p_{k}^{\prime}}}\cdot\|\theta_{k}\|_{W_{0}^{1,p_{k}}}.$
The first term in the right hand sign is uniformly bounded by a constant $c_{1}(l,\delta)$. We have also the following estimate by ($a.2$):
$\|\theta_{k}\|_{W_{0}^{1,p_{k}}}^{p_{k}}$ $\leq\int_{A_{l\sigma}}|\nabla u_{n,k}|^{p_{k}}+\int_{I_{(l+1)\sigma}^{c}}|\nabla u_{n,k}|^{p_{k}}$
$\leq\frac{1}{\nu}\int_{A_{l\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+\frac{1}{\nu}\int_{I_{(l+1)\sigma}^{c}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}.$
Taking into account the Weighted Young Inequality (see also (4.1)) $\text{with }\beta=\left(\delta p_{k}\right)^{\frac{1}{p_{k}}}$ and the boundedness of $\{\omega_{n,k}\}$ in $W^{-1,p_{k}^{\prime}}$, we obtain
$\|\omega_{n,k}\|_{W^{-1,p_{k}^{\prime}}}\cdot\|\theta_{k}\|_{W_{0}^{1,p_{k}}}$ $\leq\frac{1}{\beta^{p_{k}^{\prime}}p_{k}^{\prime}}\|\omega_{n,k}\|_{W^{-1,p_{k}^{\prime}}}^{p_{k}^{\prime}}+\frac{\beta^{p_{k}}}{p_{k}}\|\theta_{k}\|_{W_{0}^{1,p_{k}}}^{p_{k}}$
$\leq c_{2}(\delta,p_{k})+\delta\|\theta_{k}\|^{p_{k}}_{W_{0}^{1,p_{k}}}$
$\leq c_{2}(\delta,p_{k})+\frac{\delta}{\nu}\int_{A_{l\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$
$\ \ +\frac{\delta}{\nu}\int_{I_{(l+1)\sigma}^{c}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}.$
Now, we consider the left hand side in (4.6). Hypothesis ($a.3$) implies that
$\int_{\Omega}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\theta_{k}\geq 0.$
Moreover,
$\int_{\Omega}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}-2}\nabla u_{n,k}\cdot\nabla\theta_{k}$ $\geq\int_{A_{l\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$
$\quad-\delta\int_{I_{(l+1)\sigma}^{c}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}.$
Thus,
$\left(1-\frac{\delta}{\nu}\right)\int_{A_{l\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\leq\left(\delta+\frac{\delta}{\nu}\right)\int_{I_{(l+1)\sigma}^{c}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+c_{3}(l,\delta).$
Finally, from (4.5):
$\int_{\Omega_{1}^{\rho}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$ $\leq\int_{I_{l\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}\leq K_{1}^{(l)}\int_{A_{l\sigma}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K_{2}^{(l)}$
$\leq\frac{\nu\delta+\delta}{\nu-\delta}K_{1}^{(l)}\int_{I_{(l+1)\sigma}^{c}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K_{3}(l,\delta)$
$\leq\frac{\nu\delta+\delta}{\nu-\delta}K_{1}^{(l)}\int_{\Omega_{2}^{\rho}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K_{3}(\rho,\delta),$
and the assertion follows, choosing $\delta>0$ small enough. Namely, for $\delta<\nu$ such that
$0<\frac{\nu\delta+\delta}{\nu-\delta}K_{1}^{(l)}\leq\varepsilon.$
Theorem 4.4.
Let $c\in\mathbb{R}$ be a real number and let $\{u_{n}\}\subset W_{0}^{1,\boldsymbol{p}}(\Omega)$ be a $(CPS)_{c}$-sequence. Then $\{u_{n}\}$ is bounded in $W_{0}^{1,\boldsymbol{p}}(\Omega)$.
Proof.
We write $u_{n}=(u_{n,1},\dots,u_{n,d})$ and we define $\omega_{n}=(\omega_{n,1},\dots,\omega_{n,d})$ as follows
$\omega_{n,k}:=-\text{div}(A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}-2}\nabla u_{n,k})+\frac{1}{p_{k}}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}-g_{k}(x,u_{n}).$
Then
$-\|\omega_{n,k}\|_{W^{-1,p_{k}^{\prime}}}\cdot\|u_{n,k}\|_{W_{0}^{1,p_{k}}}$ $\leq\langle\omega_{n,k},u_{n,k}\rangle$
$=\int_{\Omega}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+\frac{1}{p_{k}}\int_{\Omega}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}u_{n,k}$
$\quad-\int_{\Omega}g_{k}(x,u_{n})u_{n,k}.$
We start by estimating the second integral:
$\int_{\Omega}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}u_{n,k}$ $=\int_{\Omega_{1}^{R}}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}u_{n,k}$
$\quad+\int_{\Omega_{2}^{R}}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}u_{n,k}$
$\leq CR\int_{\Omega_{1}^{R}}|\nabla u_{n,k}|^{p_{k}}+\gamma\int_{\Omega_{2}^{R}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}},$
where the set $\Omega_{1}^{R},\Omega_{2}^{R}$ are defined in Proposition 4.3.
Let $\gamma^{\prime}\in(\gamma,\mu-\overline{p})$ and $\varepsilon>0$ such that $\frac{CR\varepsilon}{\nu}\leq\gamma^{\prime}-\gamma$. According to Proposition 4.3:
$\int_{\Omega}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}u_{n,k}$ $\leq\frac{CR}{\nu}\int_{\Omega_{1}^{R}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+\gamma\int_{\Omega_{2}^{R}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}$
$\leq\left(\frac{CR\varepsilon}{\nu}+\gamma\right)\int_{\Omega_{2}^{R}}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K(R,\varepsilon)$
$\leq\gamma^{\prime}\int_{\Omega}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}+K(R,\varepsilon).$
Thus, there exists $c>0$ such that
$c\left(1+\|u_{n}\|_{W_{0}^{1,\boldsymbol{p}}}\right)$ $\geq f(u_{n})-\frac{1}{\mu}\langle\omega_{n},u_{n}\rangle=f(u_{n})-\frac{1}{\mu}\sum_{k=1}^{d}\langle\omega_{n,k},u_{n,k}\rangle$
$=\sum_{k=1}^{d}\frac{\mu-p_{k}}{p_{k}\mu}\int_{\Omega}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}-\sum_{k=1}^{d}\frac{1}{p_{k}\mu}\int_{\Omega}D_{s}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}u_{n,k}$
$\quad+\int_{\Omega}\left[\frac{1}{\mu}\sum_{k=1}^{d}g_{k}(x,u_{n})u_{n,k}-G(x,u_{n})\right]$
$\geq\sum_{k=1}^{d}\frac{\mu-p_{k}-\gamma^{\prime}}{p_{k}\mu}\int_{\Omega}A_{k}(x,u_{n,k})|\nabla u_{n,k}|^{p_{k}}-K(R,\varepsilon)$
$\geq\nu\sum_{k=1}^{d}\frac{\mu-p_{k}-\gamma^{\prime}}{p_{k}\mu}\|u_{n,k}\|_{W_{0}^{1,p_{k}}}^{p_{k}}-K(R,\varepsilon),$
which implies the boundedness of $\{u_{n}\}$ in $W_{0}^{1,\boldsymbol{p}}(\Omega)$ and the proof is complete. ∎
5. Main Theorem
Proof of Theorem 1.1.
We now proceed to verify the assumptions of Theorem 2.5.
Step 1: There exist a subspace $\mathcal{W}$ of finite codimension and $\rho,\alpha>0$ such that $f(u)\geq\alpha$ for every $u\in\partial B_{\rho}\cap\mathcal{W}$.
According to ($g.1$), and by the Mean Value Theorem there exists $\theta\in(0,1)$ such that
$|G(x,s)|\leq|\nabla G(x,\theta s)\cdot s|$ $\leq C\sum_{k=1}^{d}\left(1+\theta^{q_{k}-1}|s_{k}|^{q_{k}-1}+\sum_{j\neq k}\theta^{q_{j}\frac{q_{k}-1}{q_{k}}}|s_{j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}\right)|s_{k}|$
$\leq C\sum_{k=1}^{d}\left(|s_{k}|+|s_{k}|^{q_{k}}+\sum_{j\neq k}|s_{j}|^{q_{j}\frac{q_{k}-1}{q_{k}}}|s_{k}|\right).$
and
$|G(x,s)|\leq\kappa\sum_{k=1}^{d}\left(1+|s_{k}|^{q_{k}}\right).$
Therefore,
$f(u)$ $\geq\sum_{k=1}^{d}\frac{\nu}{p_{k}}\int_{\Omega}|\nabla u_{k}|^{p_{k}}-\sum_{k=1}^{d}\kappa\int_{\Omega}\left(1+|u_{k}|^{q_{k}}\right)$
$\geq\sum_{k=1}^{d}\frac{\nu}{p_{k}}\int_{\Omega}|\nabla u_{k}|^{p_{k}}-\kappa|\Omega|-\kappa\sum_{k=1}^{d}\int_{\Omega}|u_{k}|^{q_{k}}.$
Moreover, for every $k=1,\dots,d$ and by Hölder’s inequality, we also obtain
$\int_{\Omega}|u_{k}|^{q_{k}}\leq\left(\int_{\Omega}|u_{k}|^{p_{k}}\right)^{\frac{\tau_{k}}{p_{k}}}\cdot\left(\int_{\Omega}|u_{k}|^{p_{k}^{*}}\right)^{\frac{q_{k}-\tau_{k}}{p_{k}^{*}}},\quad q_{k}<p_{k}^{*},\ \tau_{k}=\left(\frac{1}{p_{k}}-\frac{1}{p_{k}^{*}}\right)\left(1-\frac{q_{k}}{p_{k}^{*}}\right).$
Thus,
$f(u)$ $\geq\sum_{k=1}^{d}\frac{\nu}{p_{k}}\int_{\Omega}|\nabla u_{k}|^{p_{k}}-\kappa|\Omega|-\kappa\sum_{k=1}^{d}\left(\int_{\Omega}|u_{k}|^{p_{k}}\right)^{\frac{\tau_{k}}{p_{k}}}\cdot\left(\int_{\Omega}|u_{k}|^{q_{k}}\right)^{\frac{q_{k}-\tau_{k}}{q_{k}}}.$
We consider $\varphi_{h}=(\varphi_{h,p_{1}},\dots,\varphi_{h,p_{d}})\in W_{0}^{1,\boldsymbol{p}}(\Omega)\cap L^{\infty}(\Omega;\mathbb{R}^{d})$ where $\varphi_{h,p_{k}}$ is the eigenfunction of the problem:
$\begin{cases}-\Delta_{p_{k}}\varphi_{h,p_{k}}=\lambda_{h,p_{k}}\varphi_{h,p_{k}}&\text{in $\Omega$}\\ \varphi_{h,p_{k}}=0&\text{on $\partial\Omega$},\end{cases}$
and we define ${\mathcal{W}}_{h_{0}}:=\overline{\text{span}\{\varphi_{h}\ :h\geq h_{0}\}}$. We recall that $\lambda_{h,p_{k}}\to+\infty$ for every $k=1,\dots,d$ as $h\to+\infty$, see for instance [6, 18]. Furthermore, by [6, Proposition 5.4], we have that
(5.1) $\lambda_{h_{0},p_{k}}\int_{\Omega}|u_{k}|^{p_{k}}\leq\int_{\Omega}|\nabla u_{k}|^{p},\qquad\text{for every $u=(u_{1},\dots,u_{d})\in\mathcal{W}_{h_{0}}$.}$
Hence, for every $u\in\mathcal{W}_{h_{0}}$, applying (5.1) and Sobolev’s embedding yields
$f(u)$ $\geq\sum_{k=1}^{d}\frac{\nu}{p_{k}}\int_{\Omega}|\nabla u_{k}|^{p_{k}}-\kappa\sum_{k=1}^{d}\left(\int_{\Omega}|u_{k}|^{p_{k}}\right)^{\frac{\tau_{k}}{p_{k}}}\cdot\left(\int_{\Omega}|u_{k}|^{q_{k}}\right)^{\frac{q_{k}-\tau_{k}}{q_{k}}}-\kappa|\Omega|$
$\geq\frac{\nu}{\max\{p_{1},\dots,p_{d}\}}\sum_{k=1}^{d}\|u_{k}\|_{W_{0}^{1,p_{k}}}^{p_{k}}-\kappa\sum_{k=1}^{d}\lambda_{h_{0},p_{k}}^{-\frac{\tau_{k}}{p_{k}}}\left(\int_{\Omega}|\nabla u_{k}|^{p_{k}}\right)^{\frac{\tau_{k}}{p_{k}}}\cdot\left(\int_{\Omega}|u_{k}|^{q_{k}}\right)^{\frac{q_{k}-\tau_{k}}{q_{k}}}-\kappa|\Omega|$
$\geq\frac{\nu}{\max\{p_{1},\dots,p_{d}\}}\sum_{k=1}^{d}\|u_{k}\|_{W_{0}^{1,p_{k}}}^{p_{k}}-\kappa\sum_{k=1}^{d}\lambda_{h_{0},p_{k}}^{-\frac{\tau_{k}}{p_{k}}}\|u_{k}\|_{W_{0}^{1,p_{k}}}^{q_{k}}-\kappa|\Omega|.$
Using the elementary inequality
$\left(\sum_{k=1}^{d}a_{k}\right)^{\gamma}\leq C_{d}\sum_{k=1}^{d}a_{k}^{\gamma},\qquad a_{k}\geq 0,\ \gamma\geq 1,$
where one can take $C_{d}=d^{\gamma-1}$, we obtain in particular
$\left(\sum_{k=1}^{d}a_{k}\right)^{\tilde{p}}\leq C_{d}\sum_{k=1}^{d}a_{k}^{\tilde{p}},\qquad\tilde{p}:=\min\{p_{1},\dots,p_{d}\}.$
Moreover, by Young’s inequality,
$a_{k}^{\tilde{p}}\leq\frac{\tilde{p}}{p_{k}}a_{k}^{p_{k}}+\frac{p_{k}-\tilde{p}}{p_{k}}\leq a_{k}^{p_{k}}+1,$
for every $k=1,\dots,d$. Therefore,
$\left(\sum_{k=1}^{d}a_{k}\right)^{\tilde{p}}\leq C_{d}\sum_{k=1}^{d}a_{k}^{p_{k}}+C_{d}d.$
This implies that there exist positive constants $c_{1},c_{2}>0$ (depending only on $d$ and $\tilde{p}$) such that
$\sum_{k=1}^{d}a_{k}^{p_{k}}\geq c_{1}\left(\sum_{k=1}^{d}a_{k}\right)^{\tilde{p}}-c_{2}.$
Hence, setting $a_{k}=\|u_{k}\|_{W_{0}^{1,p_{k}}}$, we conclude that
$f(u)$ $\geq\frac{\nu}{\max\{p_{1},\dots,p_{d}\}}\sum_{k=1}^{d}\|u_{k}\|_{W_{0}^{1,p_{k}}}^{p_{k}}-\sum_{k=1}^{d}\lambda_{h_{0},p_{k}}^{-\frac{\tau_{k}}{p_{k}}}\|u_{k}\|_{W_{0}^{1,p_{k}}}^{q_{k}}-\kappa|\Omega|$
$\geq\frac{\nu}{\overline{p}}c_{1}\left(\sum_{k=1}^{d}\|u_{k}\|_{W_{0}^{1,p_{k}}}\right)^{\tilde{p}}-\left(\kappa|\Omega|+\frac{\nu}{\overline{p}}c_{2}\right)+o_{h_{0}}(1),$
where $o_{h_{0}}(1)\to 0\ \text{as $h_{0}\to+\infty$.}$ Then, $f(u)\geq\alpha>0$ on $\mathcal{W}_{h_{0}}\cap\partial B_{\rho}$ for a suitable choice of $\rho>0$ and for $h_{0}\geq 1$ large enough. Indeed, for every $u\in\partial B_{\rho}$, we have that
$f(u)\geq\frac{\nu}{\overline{p}}c_{1}\rho^{\tilde{p}}-\tilde{\kappa}+o_{h_{0}}(1)\geq 1+o_{h_{0}}(1),$
for a suitable choice of $\rho>0$. Then, $f(u)\geq\alpha$ for $h_{0}\geq 1$ large enough and $\alpha>0$.
Step 2: For every subspace $V\subset W_{0}^{1,\boldsymbol{p}}(\Omega)$ of finite dimension there exists $R>\rho$ such that $f\leq 0$ in $B_{R}^{c}\cap V$.
Let $\varphi\in V\cap L^{\infty}(\Omega;\mathbb{R}^{d})$ with $\|\varphi\|=1$. By ($g.2$) we have that
$G(x,s)\geq b_{0}(x)|s|^{\mu}-a_{0}(x),\ \ \text{with $b_{0}\in L^{1}(\Omega),\ a_{0}\in L^{1}(\Omega)$},$
see, for instance, [20, Proof of Theorem 3.1, page 20]. Therefore,
(5.2) $f(t\varphi)\leq\sum_{k=1}^{d}\frac{Ct^{p_{k}}}{p_{k}}-t^{\mu}\int_{\Omega}b_{0}(x)|\varphi|^{\mu}-\|a_{0}\|_{L^{1}}\to-\infty,$
as $t\to+\infty$.
Since $W_{0}^{1,\boldsymbol{p}}(\Omega)\cap L^{\infty}(\Omega;\mathbb{R}^{d})$ is dense in $W_{0}^{1,\boldsymbol{p}}(\Omega)$, we have that for each $v\in V$ there exists $\{v_{n}\}\subset V\cap L^{\infty}(\Omega;\mathbb{R}^{d})$ with $v_{n}\to v$ in $W_{0}^{1,\boldsymbol{p}}$. Therefore,
$f(tv)=\lim_{n\to\infty}f(tv_{n})\leq 0,$
for $t>0$ large enough by (5.2).
Step 3: The functional $f$ satisfies the $(PS)$-condition.
Let $c\in\mathbb{R}$ and let $\{u_{n}\}\subset W_{0}^{1,\boldsymbol{p}}(\Omega)$ be a $(PS)$-sequence at level $c$. According to Corollary 2.8-$(ii)$, $\{u_{n}\}$ is also a $(CPS)$-sequence and Theorems 4.4 and 4.2 imply that the $(CPS)$-condition holds. Corollary 2.8-$(iii)$ allow us to conclude that the functional satisfies also the $(PS)$-condition.
Now, according to Theorem 2.5 we find a sequence of critical points $\{u_{n}\}\subset W_{0}^{1,\boldsymbol{p}}(\Omega)$ for $f$ such that $f(u_{n})\to+\infty$ and Theorem 3.1 proves that $u_{n}\in W_{0}^{1,\boldsymbol{p}}(\Omega)\cap L^{\infty}(\Omega;\mathbb{R}^{d})$, concluding the proof. ∎